12 January, Tuesday.
It is sunny, with no wind in the morning. I cannot believe it, but it is +24 C in my tent in the morning!
Then, a moderate wind starts to blow from the South-East during the day and in the evening, it changes direction to the West.
A part of the Station people has left for the traverse to the Coast, including our guides Sanne and Alain. During a recognition visit to the ridge we have found the square marked by Steve Roberts last year in the area 18 marked by Damien Ertz in 2007. We would like to make it as a reference area. It starts right after the 2nd antenna for the GPS and finishes before the cairn on top of the ridge. I saw several microscopical slides installed by Annick last year. We have selected 2 places for OTCs. Around 11 o’clock, we went to the ridge to install the OTCs. David Gladsteen was already waiting for us with a camera. He will use the videos from a briefing. He intends to use the OTC installation for a short video that will be placed on the Web-site of IPF in a few weeks. In addition, Josef collected some gravel from various sites around OTCs (integral soil sample).
OTC 1 was installed before the cairn in the middle of the square marked by Steve last year. The temperature inside was +3.2 C, and outside, it reached +0.3 C. However, the air temperature was freezing, -7.2 C. The contrast between the temperatures of the air and the rock surfaces is amazing! The rocks can really accumulate a lot of heat during sunny days!
We have installed a temperature microsensor inside the OTC, and one outside, about 30 cm further as a control. Readings of the temperature started at 16.00 and will continue each 10 minutes. We have surrounded the area with OTC 1 with 3 bamboo poles fixed by stones between rocks. All of them were connected with a red 3 mm rope. One of bamboos have a red flag from JARE 51. It can be easily seen from the station. We'll see whether they will survive a winter. We had fantasies that a webcam could be installed at a window of the station and we dreamt of being able to control in real time if the OTC was still OK.
The OTC 2 was installed behind the cairn, near the number 19 painted by Damien in 2007. The temperature inside was +4.9 C, and outside, it was -0.3 C. We installed also a humidity sensor. All sensors started to read at 16.00 o’clock, with an interval of 10 mins. We surrounded the area with bamboo poles also.
Looking for the places where living communities could be seen (like the black biofilm above), showed that these places should be protected from the wind, exposed to the sun, below a snow deposit that will give melt water.
Thus, when we went back to the Station, we had reached important results with the installation of 2 OTCs and data loggers, and protective barriers (bamboo poles and red ropes).
Tomorrow we need to :
1) go to the ridge after 16.00 (one day of functioning), collect the data loggers where temperatures and humidity are recorded, and read the data from the loggers. If everything went well, we will install them again, so that they can function till the end of January. Before leaving Utsteinen, I will set them for a longer sampling period (one data per 3 hours, that will enable records for more than a year).
2) In the morning we'll go to the cryoconite field with a drill to test the depth and size of cryoconites. Don't forget: drill, pH-meter, Winkler test, ruler.
3) Insulate the fluorometer that allows to test the photosynthetic efficiency.
4) Develop a system that will allow to mark a 1 cm area of biofilm without disturbing it. It will be necessary for the fluorometer observations (idea of Josef)
5) Test of the fluorometer in a field conditions with a good weather.
The weather forecast is good for next days, but starting from Friday, the weather will become worse. It will be bad on Saturday and may be Sunday. We need to find what to do during these days, probably we'll not be able to go to Pingvinane. In Antarctica, the weather has the final word about a lot of things.